OpenStudy (anonymous):

A porch has the shape of a quarter circle with a diameter of 5 ft. What is the approximate area of the porch, rounded to the nearest square foot? Use 3.14 to approximate for pi .

OpenStudy (anonymous):

I got done with this one. Can you help me with another @iGreen

OpenStudy (igreen):

Act like it's a whole circle, find the area, then divide by 4. A = pi * r^2 A = 3.14 * 2.5^2 Multiply 3.14 * 2.5 * 2.5

OpenStudy (anonymous):

Which figures can be used to find the area of the composite figure? Choose exactly two answers that are correct. A. a rectangle with an area of 34 cm2 B. a rectangle with an area of 72 cm2 C. a quarter circle with an area of cm2 D. a quarter circle with an area of cm2

OpenStudy (anonymous):

theres the pic

OpenStudy (anonymous):

hello

OpenStudy (igreen):

Act like it's a whole circle, find the area, then divide by 4. A = pi * r^2 A = 3.14 * 2.5^2 Multiply 3.14 * 2.5 * 2.5

OpenStudy (anonymous):

are you there

OpenStudy (anonymous):

@almightydmc23 @1351250faith @Jhannybean @jl700

OpenStudy (igreen):

Can you multiply 3.14 * 2.5 * 2.5?

OpenStudy (anonymous):

yeah

OpenStudy (igreen):

What do you get?

OpenStudy (anonymous):

19.625

OpenStudy (igreen):

Yes, now divide by 4.

OpenStudy (anonymous):

ok. 78.5

OpenStudy (igreen):

Divide, not multiply

OpenStudy (anonymous):

i want to get a second opinion just to make sure ur right

OpenStudy (anonymous):

i did divide by 4

OpenStudy (igreen):

\(19.625 \div 4 \ne 78.5\)

OpenStudy (anonymous):

4.90

OpenStudy (igreen):

Yes, so that's our answer.

OpenStudy (mathstudent55):

You are still solving the first problem? I thought you said you already had done it.

OpenStudy (igreen):

Oh..I just noticed that..sorry..

OpenStudy (anonymous):

I did. he wouldnt do the second if he didnt help me with the fist one

OpenStudy (anonymous):

what about the second

OpenStudy (mathstudent55):

Ok.

OpenStudy (igreen):

The rectangle has length of 9 and a width of 8, multiply 9 * 8

OpenStudy (anonymous):

72

OpenStudy (anonymous):

ok then what

OpenStudy (igreen):

Yes, so the rectangle has an area of 72, now for the quarter circle. It has a radius of 8: A = pi * r^2 A = 3.14 * 8^2 Simplify

OpenStudy (anonymous):

50.24

OpenStudy (mathstudent55):